1) The U.S. Treasury issued a 7-year maturity, $1000 par value bond exactly 3 years ago. The bond pays a nominal coupon rate of 12%. The coupon payments are paid semi-annually The most recent coupon payment (the sixth coupon payment) was made yesterday. Your required rate of return from the bond is 10% per year What is the price of the bond today? If the bond can be called in 5 years from the date of issue at a call price of $1120, and you bought the bond today at the price computed in part (a), what would be your yield-to-call (YTC) from the bond? a) b) 2) Scan Solutions, Inc. issued a 15 year maturity, 6% semi-annual coupon paying bond 7 years ago. You purchased the bond at par value at the time of issue. You intend to sell the bond now. Similar maturity, similar risk bonds currently yield 8.2% per year What price do you expect to receive for the Scan Solutions bond if you sell it today? What is your realized yield if you succeed in selling the bond today at a price of $870? At a price of $885? a) b)

Please show the steps to finding the answer using a *Financial Calculator*! Thank you.
Answer #1

1)

A B C D E F G H I J K L
2 a)
3 Par value (F) $1,000
4 Interest rate (Coupon rate) 12.00%
5 Market demanded return 10.00%
6 Time to maturity 4 Years
7
8 Interest is paid twice a year i.e. semiannual.
9 Semiannual coupon (C) $60.00 =D3*D4/2
10 Semiannual Period (n) 8 =D6*2
11 Semiannual YTM (i) 5.00% =D5/2
12 Price of bond is the present value of future cash flows.
13 Price of Bond can be found by entering the following values in financial calculator:
14 FV $1,000
15 PMT $60.00
16 I/Y 5.00% (Enter 5 in place of 5%)
17 N 8
18
19 After entering the above values, press PV which will given the bond price in negative.
20
21 Bond Price ($1,064.63) =PV(D16,D17,D15,D14)
22
23 Hence Bond Price is $1,064.63
24
25 b)
26
27 Par value (F) $1,000
28 Interest rate (Coupon rate) 12.00%
29 Call Price $1,120
30 Time to call 2 Years
31 Current Price $1,064.63
32 Interest is paid twice a year i.e. semiannual.
33 Semiannual coupon (C) $60.00 =D27*D28/2
34 Semiannual Period (n) 4 =D30*2
35
36 In the case bound will be bought at current price and coupon will be received for 2 years and after that bond will be called at $1,120.
37 Yield to call can be found by entering the following values in financial calculator:
38 FV $1,120
39 PMT $60.00
40 PV ($1,064.63) (Current Price is cash outflow)
41 N 4
42
43 After entering the above values, press I/Y which will give the yield to call.
44
45 Yield to call 6.81% =RATE(D41,D39,D40,D38)
46
47 Hence Yield to call is 6.81%
48